Displaying reports 1041-1060 of 1079.Go to page Start 49 50 51 52 53 End
Reports until 11:36, Wednesday 16 August 2023
Controls Optimal Controls
Ian MacMillan - posted 11:36, Wednesday 16 August 2023 - last comment - 11:36, Wednesday 16 August 2023(11207)
Correcting the Flat FOM gain to always be 1

[Ian, Lee]

NOTE: this should have been posted on June 21 2023 as this is the date that it was writen, but it was in my drafts for some reason. I have added latex since then

In the H2 system the gain for the flat FOM (F2) is 1/F1_gain, where F1_gain is the gain of the BNS FOM (F1). This works fine for the H2 norm because that is made up of the composite RMS from both FOMs so as one FOM's gain increases and the others decreases they cancel out. However, this is not the case in the H_infinity case. The H_infinity norm is the maximum value for the frequency response from the Zinf input to the Flat FOM output this value must be less than or equal to the gamma that we are using. Currently we are searching for a gamma around 1. The equation for the frequency response along the  Zinf input to the Flat FOM output is:

\[ \left | \frac{K(\omega)P(\omega)}{1-K(\omega)P(\omega)} F_{\text{flat}}(\omega)\right |\leq \gamma \]

where \( K(\omega)\) is the controller response, \(P(\omega)\) is the plant response and \(F_{\text{flat}}(\omega)\) is the Flat FOM's response. If the environmental noise is large like at low frequencies then \( K(\omega)P(\omega) \) must be large to control it. Thus at low frequencies the above equation becomes,

\[ \lim_{KP \to \infty }\left | \frac{K(\omega)P(\omega)}{1-K(\omega)P(\omega)} F_{\text{flat}}(\omega)\right | \approx\left |F_{\text{flat}}\right | \leq \gamma \]

Thus if we use the F2_gain=1/F1_gain then as F1_gain gets very small F2_gain gets very large. When F2_gain is very large, the gamma we are searching for (around 1) is much less than the actual H_infinty norm which is on the order of the F2_gain and the solver fails. This is what was causing the problems with the RMS plot. 

We switched the F2 gain to be all ones and acquired attached plot.

Note: when this PDF is printed the second color bar has the same colors as the first color bar.


Improvements for code and plots for paper:

  1. Fix the color bar problem (noted above) for printing.
  2. Make the color bar discreet
  3. decrease the number of F1_gain points so that the plot is not so crowded 
  4. Maybe add an H2 boundary line. See how it looks
  5. change the F1_gain color bar. maybe connect like F1_gain points with a subtle line.
  6. Add an example of the model maker (getSPOFF) to Buzz
  7. make the code transpose the model right before solving in the H_infinity case
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 10:54, Friday 23 June 2023 (11210)

Infinite tweaks made.

Seeking control perfection.

Efficiency thrives.

Electronics General
Ian MacMillan - posted 16:38, Wednesday 28 June 2023 - last comment - 16:51, Wednesday 28 June 2023(11222)
Set Up Temporary Soldering Station

[Ian, Torrey]

We set up a temporary soldering station in the swing space. this is just until the full lab ins built then we can move it over there. It is also so that we could move all of the electronics equipment off of the optical table. it is very crowded right now so we should buy some storage for all of the stuff to clean up the area a little bit, but for now at least it is off the optics table. We also need to order some solder because we don't have any.

See picture

Images attached to this report
Comments related to this report
Torrey Cullen - 16:51, Wednesday 28 June 2023 (11223)

Melting metal joins,

Solder's fiery touch binds.

Alex buy solder.

Lab Infrastructure Networking
Ian MacMillan - posted 14:48, Wednesday 28 June 2023 - last comment - 16:10, Wednesday 28 June 2023(11219)
Added Network to Swing Space

I set up the network for in the swing space. Both the 2.4GHz and 5GHz networks combined into one network. We can turn off the 2.4GHz network if it starts to be a problem. 

Router model: ASUS - AX6000 Dual Band Wi-Fi 6 Router, Model: RT-AX88U Pro

Network Name: pLANk time

Network and Admin Passwords: LIGO Secrets or written on router

Admin site:  http://192.168.50.1

Comments related to this report
Torrey Cullen - 16:10, Wednesday 28 June 2023 (11220)

In swing space's realm,

Ian supplies internet.

LIGO's secrets held.

Lasers General
Torrey Cullen - posted 14:48, Wednesday 28 June 2023 - last comment - 16:25, Wednesday 28 June 2023(11218)
cryo lab update and unknown oscillations request

The Mach–Zehnder in the cryo lab was installed with a beamsplitter mounted on a 3 port mount. Went to install the second readout of the Mach–Zehnder and discovered this. I have since reinstalled the beamsplitter with an appropriate mount, put the other 1811 PD with a lens in front of it on the second readout path, and recovered the alignment on both paths. Minor alignment is still needed to maximize visibility but there are fringes.

 

Additionally, the source of the oscillations on the PD in reflection of the cavity are still unresolved. Here are some things I know about it:

1) I don't believe they are physical power fluctuations with the laser. As you can see from the screen shot, those fluctuations are between ~3-4 volts on the PD. If these were physical power fluctuations they would show up on a power meter; they don't.

2) They don't seem to be a saturation issue as you can use the half wave plate + PBS to control the amount of power to the cavity and they still occur at very low power.

3) Thought maybe an electrical issue. I substituted the power source for a new 15V power source that we just bought from Newport. Aaron suggested having the moku and PD power from the same power strip so they have a common ground. Also I've tried plugging in the PD to different outlets all around the room. Nothing has worked.

4) The new 1811 I installed in the Mach-Zehnder path (that doesn't see the cavity at all) does not have these fluctuations.

My only other idea is maybe the suspension of this chamber that the cavity is housed in (if it has a suspension) is causing it but I'm unsure how to mitigate/test for that. If anyone has any ideas, please let me know.

Images attached to this report
Comments related to this report
Ian MacMillan - 16:25, Wednesday 28 June 2023 (11221)

Unresolved ripples

in the Mach-Zender's dance,

knowledge to unfold

 

Alternate 7-7 Poem:

One port, two port, three port, four,

Dr. Torrey just wants more.

Controls Optimal Controls
Ian MacMillan - posted 12:16, Wednesday 28 June 2023 - last comment - 11:33, Wednesday 05 July 2023(11217)
Using non-stabalized ASC model in Buzz

In all previous tests I have been using a model of the ASC system that was fit from data using IIRRational in the 'ASC' Folder of the 300m repo. The ZPK representation (See attachment T_ASC_CHARD_P_Unstable.yml) that it produced had right half plane poles meaning the filter was unstable. We then stabilized this by hand by flipping the right hand plane poles into left hand plane poles, i.e. 4.9234+12442i  -> -4.9234+12442i (See attachment T_ASC_CHARD_P.yml). We then used the stabilized version in all of our testing. 

I went back and tried to see if the solver could handle the the unstable ASC model. I added the model  (See attachment ASC_BH2_solver2_unstable.mat) to the buzz 'ExampleModels' folder in the buzz repo and tried it. The solver fails when calculating:

'Q = solve_continuous_are_scipy(a=A.T, b=C.T, q=V1i, r=bh.D2 @ bh.D2.T, e=None, s=V12i, balanced=True)'

giving the error:

'LinAlgError: Failed to find a finite solution.'

 

Non-image files attached to this report
Comments related to this report
Ian MacMillan - 11:33, Wednesday 05 July 2023 (11230)

See [11229] for the fix

GQuEST General
Chris Stoughton - posted 07:31, Tuesday 27 June 2023 - last comment - 12:13, Tuesday 27 June 2023(11213)
GQuEST in One Page
Here is a one-page summary of GQuEST. Let's discuss!
Non-image files attached to this report
Comments related to this report
Ian MacMillan - 12:13, Tuesday 27 June 2023 (11216)

A tribe called GQuEST,

summary is one page long,

elevator pitch.

GQuEST Output Filter Cavities
Daniel Grass - posted 15:54, Monday 26 June 2023 - last comment - 12:11, Tuesday 27 June 2023(11183)
Bowtie Cavity Mechanical Assembly

I added helicoils to a bowtie cavity, the piezo bases, and the mirror deformation bases. I also tapped the piezo bases and the mirror deformation bases with SM1 threads. I fully assembled a bowtie cavity as a fit check. Future steps: Add helicoils to the remaining bowtie cavities. Test the static mirror deformation. Purchasing required for bowtie cavities: Indium wire/foil 1 in Mirrors Piezos are currently being shipped 06/06/2023 Added helicoils to remaining bowtie cavities For some reason, one of the piezo bases had 3 through holes with too small a diameter. I reached out to Hubs to get a refund on that part. I can drill out the holes by ~4 thou so that the screws fit. Other than this issue, I was able to successfully fit check everything. I also successfully stacked two bowties on top of each other.

Comments related to this report
Ian MacMillan - 12:11, Tuesday 27 June 2023 (11215)

Helicoils on,

cavities stacked high like cash,

thou is too cool, say thousandth 

Equipment General
Daniel Grass - posted 15:53, Monday 26 June 2023 - last comment - 12:07, Tuesday 27 June 2023(11212)
Holometer Disassembly June 20-June 23
While at Fermilab, I disassembled a lot of parts like the central vessels, two end cubes, and a lot of gate values and bellows. I specifically prepared two central vessels and two end cubes to be shipped by emptying their contents and covering the open ports with plastic and duct tape. There are also a lot of other parts nearly ready to be shipped, like 1 foot T pipes, that will be useful for the 0.5 m interferometer. Two (b)end cubes in the shed also need to be disassembled. We will probably need a lot of small parts. The only expensive part we need, as far as I can tell, is another end cube. The remaining two cubes are being used by other experiments. Remaining work: Get central vessels (2) and end cubes (2) out of building Disassemble (b)end cubes (2) Track down remaining end cube Prepare various pipe lengths to be shipped Prepare 8 inch gate valves (4?) to be shipped Prepare bellows to be shipped Have carpenters build structure to protect everything in shipping
Comments related to this report
Ian MacMillan - 12:07, Tuesday 27 June 2023 (11214)

Disassembled parts,

Shipped vessels and cubes with care,

Labbook's tale unfolds.

Lab Infrastructure General
Ian MacMillan - posted 17:03, Friday 23 June 2023 (11211)
Table Leveling and Condition

[Ian, Torrey]

The long table on the North wall of the transition lab was leveled by putting shims under the table legs (See Table_Before_Shims.jpeg). We found that there are adjustable feet under the legs and we extended them (See Table_After.jpeg).

While extending the legs we found that the table top was not attached to the frame of the table and almost impossible to level. We tried to level it the best we could but the table in general doesnt look to be in great shape.

Controls Optimal Controls
Ian MacMillan - posted 12:59, Tuesday 20 June 2023 (11205)
Implementation of Balancing in State-Space Systems

[Ian, Lee]

Balancing state space systems is essential for getting the RMS plots correct for the comparison of controllers. When using the solution for the Lyapunov equation the solver benefits greatly. Adding the wrapped version of Slycot's AB09ND was the key to getting the H2 Lyapunov solver to correctly display the shape of the RMS H2 optimal curve. AB09ND "compute[s] a reduced order model (Ar,Br,Cr,Dr) for an original state-space representation (A,B,C,D) by using either the square-root or the balancing-free square-root Singular Perturbation Approximation (SPA) model reduction method for the ALPHA-stable part of the system" (from documentation).

Another way that was shown to be effective at helping the Laub method to work in some updates for GWINC is the TB01ID. We were wondering if this would be as effective on our code given that AB09ND is calling TB01ID. We ran the code with balancing from AB09ND, balancing from TB01ID, and no balancing.

We found that the only balancing option that gave us a segment of continuous RMS gradient was AB09ND. Both TB01ID and the no balancing showed a scatter which should not be the case.

We have attached the RMS plots for all three situations.

Non-image files attached to this report
GQuEST Meeting (GQuEST)
Chris Stoughton - posted 15:54, Monday 19 June 2023 - last comment - 15:37, Wednesday 21 June 2023(11203)
Trivial demonstration of counting statistics
This is mostly a test to see how to use jupyter notebooks for communications. They can't at the moment, but here it is exported as a pdf.
Non-image files attached to this report
Comments related to this report
Ian MacMillan - 10:05, Tuesday 20 June 2023 (11204)

I added .ipynb to the list of allowable files

Ian MacMillan - 15:37, Wednesday 21 June 2023 (11208)

Testing uploading .ipynb file for the above code

Non-image files attached to this comment
Lasers General
Torrey Cullen - posted 11:43, Wednesday 14 June 2023 - last comment - 14:42, Thursday 22 June 2023(11196)
First acquired lock on the cryo lab west cavity

I successfully locked the cryo west cavity using the laser lock box on the moku yesterday. The main catalyst to this was that the PD being used to measure the REFL signal had a bandwidth of ~10 MHz and was attenuating the 32.7 MHz modulation signal. We ordered some newport 1811 PDs (BW = 125 MHz), and the modulation signal shot up by several orders of magnitude. Note that there is now a lens in front of the REFL PD because the diode on the 1811 is very small, also the damage threshold on these is MUCH LOWER than some of the thorlab equivalents. Don't send > 1mW towards the cavity.

First, the settings of the laser lock box that were used for initial lock can be seen in the attached text file, as well as a screen shot of the lockbox while locked. The process I was using was:

1) Make sure the TEC setpoint is roughly centered to ensure maximum range on the slow loop. Manual engage the slow loop.

2) Slide the slow loop offset until its roughly near the cavity resonance, i.e. you see the large dip in REFL PD.

3) Engage the laser lock box assist function, which is this button. You should now see the REFL PD dip is centered on the error signal zero crossing point, along with some cross-hairs highlighting the zero crossing point on the error signal. These are clickable. Click the zero point that corresponds to the RELF PD dip (sometimes the moku adds additional cross-hairs that are garbage). I'm realizing I don't have a screenshot of this, I'll grab one and comment on this post with it later today.

And thats it! Pretty easy with the moku once everything is set up. I kept the cavity locked for 10+ minutes while tapping the table, the lock seems fairly robust. A minor nuisance, in the other attached photo there seems to be some random low frequency oscillations on the REFL PD that I'm unsure where its coming from. 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 13:23, Friday 16 June 2023 (11200)

Example of zero point crossing when using the lock assist function.

Images attached to this comment
Ian MacMillan - 15:14, Friday 16 June 2023 (11202)

Laser's steady hum,

Cryo cavity now locked,

Early victory.

Torrey Cullen - 14:42, Thursday 22 June 2023 (11209)

Diagram of current setup when lock was achieved. 

Images attached to this comment
Non-image files attached to this comment
Displaying reports 1041-1060 of 1079.Go to page Start 49 50 51 52 53 End